OpenStudy (amber15):

A private plane traveled from Seattle to a rugged wilderness at an average speed of 315mph. on the return trip the average speed was 252. if the total traveling time was 6 hours, how far is Seattle to wilderness

OpenStudy (perl):



OpenStudy (amber15):

ok

OpenStudy (amber15):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

from Seattle to the wilderness d = r*t d = 315*t t = d/315 -------------------------- from the wilderness to Seattle d = r*t d = 252*t t = d/252 -------------------------- Add up the two time values and set that total equal to 6. This is valid because it takes a total of 6 hours to do the round trip (time1) + (time2) = total time d/315 + d/252 = 6 do you know how to solve for d from here?

OpenStudy (amber15):

would it be 252 for d

jimthompson5910 (jim_thompson5910):

d is unknown at this point 252 is the speed coming back

OpenStudy (amber15):

oh okay

jimthompson5910 (jim_thompson5910):

d = 252 is not the solution

OpenStudy (amber15):

would D be 840

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (amber15):

awesome your good